Carbon dioxide, which is recognized as the major contributor to global warming as a "greenhouse gas," is formed when fossil fuels are combusted, as in electrical power plants fueled by coal, oil, or natural gas. One potential way to reduce the amount of CO2 added to the atmosphere is to store it as a compressed gas in underground formations. Consider a 1000-megawatt coal-fired power plant that produces about 7×106 tons of CO2 per year.

(1) Assuming ideal gas behavior, 1.00 atm, and 17 ∘C, calculate the volume of CO2 produced by this power plant.

(2) If it is stored underground as a gas at 33 ∘C and 90 atm , what volume does it occupy?
1
Expert's answer
2017-10-23T08:55:41-0400
n(CO 2 ) = m/M = 7*10 6 / (12+16*2) = 1.59*10 5
pV = nRT
a) V(CO 2 ) = 1.59*10 5 * 8.314*290 / 101.325 = 3.78 * 10 6 (L)
b) V(CO 2 ) = 1.59*10 5 * 8.314*306 / (101.325*90) = 4.43 * 10 4 (L)
Answer
a) V(CO 2 ) = 3.78 * 10 6 L
b) V(CO 2 ) = 4.43 * 10 4 L